OpenStudy (anonymous):

Y''+9y'=t^2 e^(3t) + 6 solved via y(t)=yc+yp?

OpenStudy (anonymous):

homogenous: m^2+9m=0 m(m+9)=0 m = 0 and m = -9 Yh(x) = C1 + C2 e^-9x y1 = e^0 = 1 y2 = e^-9x wronskian = W[y1,y2] = 1 e^-9x 0 -9 e^-9x = -9e^-9x - 0 = -9 e^-9x not equal to 0, means linear free continue......

OpenStudy (anonymous):

yes